News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 14 September 2001 10.00am EDT

During the last week, Chandra completed the observing schedule as planned.

A real-time procedure was completed on Sep 12 to update the on-board ephemeris.

The schedule of observations for the coming week is shown below and includes a series of 9 calibration observations of the dark moon and 4 observations of the Comet 19P/Borrelly. The observation of EV Lac is coordinated with the Hubble Space Telescope and the Very Large Array.

------------------------------------------
RX J0852.0-4622         ACIS-S       Sep 16
Tau Ceti                ACIS-S       Sep 17
PSS 0059+0003           ACIS-S
BRI 0103+0032           ACIS-S
3C120                   ACIS-S       Sep 18
Radiation Belts
NGC3147                 ACIS-S
WZ SGE                  ACIS-S
CAS-A                   HRC-S        Sep 19
EV Lac                  ACIS-S/HETG
Radiation Belts                      Sep 20
1WGA_J1233.3+6910       ACIS-S       Sep 21
MWC297                  ACIS-I
SAX J1752.3-3138        ACIS-S       Sep 22
Dark Moon (9 obs)       ACIS-I
Comet 19P/Borrelly (4)  ACIS-S
Puppis N Knot           ACIS-S       Sep 23
1RXS J1708-4009         HRC-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports